OpenStudy (anonymous):

10x+30(-10)=10+30(-10)

OpenStudy (anonymous):

\[10x + 30(-10)=10+30(-10)\]

OpenStudy (anonymous):

\[10x-300=10-300\]

OpenStudy (anonymous):

10x = 10 x= 1

OpenStudy (anonymous):

Ah i see. Thanks.

OpenStudy (anonymous):

You're welcome